# Structure-Fair Quantum Circuit Complexity: An Information-Theoretic Lower Bound

### v1

@@ Abstract
Fairly quantifying the complexity of quantum states that possess intrinsic structures (such as symmetries or encodings) is a fundamental challenge for benchmarking quantum technologies. We introduce the "Reference-Contingent Complexity" (RCC). Its core idea is to utilize the quantum relative entropy to measure the extent to which a state deviates from its "structured vacuum" (the maximum entropy state within its constrained subspace), thereby only pricing the generation of non-trivial information. We establish a central theorem, establishing that the RCC is a rigorous information-theoretic lower bound on universal quantum circuit complexity, a bound that features a linear leading term, a universal logarithmic correction, and a precise physical correction for the non-uniformity of the spectral distribution. Furthermore, we establish an operational protocol based on quantum hypothesis testing, which connects this theoretical lower bound to experimentally accessible observables. This work provides a structurally-fair yardstick for quantum technologies and offers a new perspective for exploring the intrinsic connection between computational cost and the generation of non-trivial information.
@@ 2.3 Definition of Reference-Contingent Complexity
Eq. (2.3) reveals the physical meaning of RCC. Its numerator term, $(\log d_{R}-S(\rho))$, represents the “information deficit” or “degree of order” of the quantum state $\rho$ within its respective structured subspace $H_{R}$. $\log d_{R}$ is the maximum Shannon information capacity of this subspace (i.e., $S(\sigma_{R})$), while $S(\rho)$ is the von Neumann entropy actually contained in the state $\rho$. The difference between the two is precisely the degree to which the state deviates from “complete disorder,” and it is also precisely the amount of non-trivial information that must be “injected” by any quantum circuit that prepares the state. At the same time, the denominator term $\log\Gamma_{R}$ provides a precise calibration of the fundamental unit of computational resources. Therefore, the physical picture of RCC is: the structurally fair complexity of a quantum state is equal to the minimum amount of information required to be injected for its preparation, divided by the maximum information bandwidth that a single complete computational operation can provide. To specifically illustrate how RCC achieves structural fairness, we once again consider a stabilizer code that encodes $k$ logical qubits into $N$ physical qubits. The dimension of its code subspace $H_{R}$ is $d_{R}=2^{k}$. For a logical pure state $\rho_{L}$ within the code space, its von Neumann entropy is $S(\rho_{L})=0$. Assuming logical gates can act arbitrarily on the $k$ logical qubits with a universal gate set of cardinality $g_{L}$, and that a single operation has an addressable range covering all $k$ logical qubits, i.e., $|S_{R}|=k$. According to Eq. (2.3) and the definition $\Gamma_{R}=g_{L}\cdot|S_{R}|$, its RCC is $C_{R}(\rho_{L})=k/\log(g_{L}\cdot k)$. This result aligns with physical intuition: the complexity of this logical pure state is proportional to the number of logical information bits it carries, $k$, and is independent of the number of physical qubits used for encoding, $N$, while the denominator correctly reflects the cost of the logical operation set. This clearly demonstrates how RCC, by its very definition, automatically strips away the ‘free’ encoding structure and only prices the non-trivial logical information. In Appendix C, we provide more detailed implementation guidance by treating this example alongside another core physical scenario (fixed particle number sector) as canonical case studies.
@@ 2.4 A Note on Units: The Structon
The physical interpretation of this unit is: against the background of a given “structured vacuum” $\sigma_{R}$, 1 $\mathrm{st}_{R}$ represents the complexity required to generate a non-trivial structural unit equivalent to the information of a $\Gamma_{R}$-alphabet”. As an independent and self-consistent notational convention, the introduction of $\mathrm{st}_{R}$ does not alter the form or proof of any theorem in this paper; all correction terms in the main theorem, Theorem 3.1 (spectral skew and logarithmic overhead), when measured in bits or nats, are naturally converted to the $\mathrm{st}_{R}$ unit via Eq. (2.5). To maintain compatibility with traditional information-theoretic expressions, the main text will primarily use the dimensionless form $D/\log\Gamma_{R}$; readers may directly interpret this as a physical quantity in units of $\mathrm{st}_{R}$. To ensure clarity and reproducibility, we recommend following a minimal set of guidelines when reporting results using $\mathrm{st}_{R}$:
@@ 6.3 Resolving the Paradox of Pure State Complexity through an Observational Window
Proof Outline. Let $\sigma=E_{\Xi}(\rho)$. Apply a pinching map $\Delta_{P}(\sigma)=\sum_{i}P_{i}\sigma P_{i}$ to $\sigma$. By the entropy-increasing property of pinching, $S(\sigma)\leq S(\Delta_{P}(\sigma))$. At the same time, the entropy of the pinched state satisfies the block-entropy decomposition $S(\Delta_{P}(\sigma))=H(\{p_{\Xi,i}\})+\sum_{i}p_{\Xi,i}S(\sigma_{i})\leq H(\{p_{\Xi,i}\})+\sum_{i}p_{\Xi,i}\ln r_{i}$, where $\sigma_{i}=P_{i}\sigma P_{i}/p_{\Xi,i}$. Substituting this chain of inequalities back into the definition of $C_{R,\Xi}(\rho)$, $(\ln d_{R}-S(\sigma))/\ln\Gamma_{R}$, yields Eq. (6.3.1). (For technical details on the entropy increase due to pinching and the block-entropy decomposition used in this proposition, see Ref. [22]; for the rank-one dephasing entropy bound and experimental protocol, see D; for the axiomatic background of windowed monotonicity and notational/unit conventions, see Appendix I.) The physical insight of Eq. (6.3.1) is particularly crucial: it provides an operational "lever" to break the complexity degeneracy of pure states under a finite observational capability. For any pure state, the upper bound of the leading term of its ideal global complexity $C_{R}(\rho)$ is $\ln(d_{R})/\ln(\Gamma_{R})$. However, through a fixed local observational window $E_{\Xi}$, the value of $S(E_{\Xi}(\rho))$ for different pure states will differ significantly due to their different entanglement structures, thus creating a gap in their $C_{R,\Xi}(\rho)$ values. The significance of Eq. (6.3.2) is that it transforms $C_{R,\Xi}(\rho)$ from a theoretical construct into a physical quantity with direct operational meaning. While the definition of $C_{R,\Xi}(\rho)$ itself already conceptually solves the pure state degeneracy problem through $S(E_{\Xi}(\rho))$, Eq. (6.3.2) provides, from a more practical direction, an operational lower bound determined entirely by the classical measurement outcomes $\{p_{\Xi,i}\}$. This forms a key bridge connecting abstract theory with laboratory reality and gives rise to a powerful scheme for actively probing the latent complexity within pure states, which we term the "Experimental Squeeze of the Complexity Spectrum." Its core logic is that the choice of the measurement basis $\{P_{i}\}$ reflects the experimenter’s "informed probing" of the system’s intrinsic structural complexity. For a highly structured pure state, such as a graph state defined by a stabilizer group $S$, if we "intelligently" choose a measurement basis that is "aligned" with its stabilizer generators, then according to stabilizer theory, the measurement outcomes will be highly deterministic ($\pm 1$). This will cause the probability distribution $\{p_{\Xi,i}\}$ to become extremely sparse, with its Shannon entropy $H$ approaching zero. According to Eq. (6.3.2), this will cause its measurable complexity lower bound $C_{R,\Xi}(\rho)$ to be "squeezed" upwards, approaching its theoretical upper limit $\ln(d_{R})/\ln(\Gamma_{R})$. Conversely, a simple product state, lacking intrinsic non-local correlations, will inevitably appear more random when probed with the same complex measurement basis, resulting in a larger $H$ and a lower bound. This sharp contrast reveals a profound principle: the structural complexity of a state can only be most fully revealed when it is probed by an observer who "understands" its structural complexity. The significance of this "experimental squeeze" scheme is twofold. First, it constructs a feasible verification path for testing the core open question of this paper, Conjecture 6.1 (the near-tightness of RCC). By finding the optimal measurement bases that maximize the complexity lower bound for different classes of pure states (such as graph states, tensor network states, etc.), we can systematically establish "certificates of tightness" for their complexity. Second, it elevates the RCC framework from a static spectral measure to a complete, observation-dependent dynamical theoretical framework, in which the operational process of "finding the optimal measurement basis" is itself equivalent to the core task of "decoding the structural complexity of the quantum state."

### v2

# Structure-Fair Quantum Circuit Complexity: An Auditable Information-Theoretic Lower Bound

@@ Abstract
Quantifying the complexity of quantum states that possess intrinsic structure, such as symmetry or encoding, in a fair manner constitutes a core challenge in the benchmarking of quantum technologies. This paper introduces the Reference-Contingent Complexity (RCC), an information-theoretic measure calibrated by the available quantum operations. The core idea is to leverage the quantum relative entropy to quantify the deviation of a quantum state from its "structured vacuum"—namely, the maximum entropy state within its constrained subspace—thereby only pricing the process of creating non-trivial information. Our central result is a key theorem that rigorously proves the RCC serves as a lower bound for the complexity of any universal quantum circuit. This lower bound is comprised of a linear dominant term, a universal logarithmic correction, and a precise physical correction term that accounts for non-uniformity in the spectral distribution. Crucially, we establish a set of operational protocols, grounded in tasks like quantum hypothesis testing, which make this theoretical lower bound experimentally "auditable." This work provides a "ruler" for quantum technology that is structure-fair and enables cross-platform comparison, thereby establishing a strictly verifiable constraint between the computational cost of the process and the structured information of the final state.
@@ 2.3 Definition of Reference-Contingent Complexity
Eq. (2.3) reveals the physical meaning of RCC. Its numerator term, $(\log d_{R}-S(\rho))$, represents the “information deficit” or “degree of order” of the quantum state $\rho$ within its respective structured subspace $H_{R}$. $\log d_{R}$ is the maximum Shannon information capacity of this subspace (i.e., $S(\sigma_{R})$), while $S(\rho)$ is the von Neumann entropy actually contained in the state $\rho$. The difference between the two is precisely the degree to which the state deviates from “complete disorder,” and it is also precisely the amount of non-trivial information that must be “injected” by any quantum circuit that prepares the state. At the same time, the denominator term $\log\Gamma_{R}$ provides a precise calibration of the fundamental unit of computational resources. Therefore, the physical picture of RCC is: the structurally fair complexity of a quantum state is equal to the minimum amount of information required to be injected for its preparation, divided by the maximum information bandwidth that a single complete computational operation can provide. To specifically illustrate how RCC achieves structural fairness, we once again consider a stabilizer code that encodes $k$ logical qubits into $N$ physical qubits. The dimension of its code subspace $H_{R}$ is $d_{R}=2^{k}$. For a logical pure state $\rho_{L}$ within the code space, its von Neumann entropy is $S(\rho_{L})=0$. Assuming logical gates can act arbitrarily on the $k$ logical qubits with a universal gate set of cardinality $g_{L}$, and that a single operation has an addressable range covering all $k$ logical qubits, i.e., $|S_{R}|=k$. According to Eq. (2.3) and the definition $\Gamma_{R}=g_{L}\cdot|S_{R}|$, its RCC is $C_{R}(\rho_{L})=k/\log(g_{L}\cdot k)$. This result aligns with physical intuition: the complexity of this logical pure state is proportional to the number of logical information bits it carries, $k$, and is independent of the number of physical qubits used for encoding, $N$, while the denominator correctly reflects the cost of the logical operation set. This clearly demonstrates how RCC, by its very definition, automatically strips away the ‘free’ encoding structure and only prices the non-trivial logical information. In Appendix C, we provide more detailed implementation guidance by treating this example alongside another core physical scenario (fixed particle number sector) as canonical case studies.It is noteworthy that, from this global perspective, all pure states ($S(\rho)=0$) within the subspace $H_{R}$ share the same maximal leading term $\log(d_{R})/\log(\Gamma_{R})$. This degeneracy is a direct consequence of the information-theoretic nature of RCC. We will revisit this profound feature in Sec. 6.3, where we demonstrate how this degeneracy is lifted by introducing a physically motivated “observational window,” thereby distinguishing the rich internal structures of different pure states.
@@ 2.4 A Note on Units: The Structon
The physical interpretation of this unit is: against the background of a given “structured vacuum” $\sigma_{R}$, 1 $\mathrm{st}_{R}$ represents the complexity required to generate a non-trivial structural unit equivalent to the information of a $\Gamma_{R}$-alphabet”. As an independent and self-consistent notational convention, the introduction of $\mathrm{st}_{R}$ does not alter the form or proof of any theorem in this paper; all correction terms in the main theorem, Theorem 3.1 (spectral skew and logarithmic overhead), when measured in bits or nats, are naturally converted to the $\mathrm{st}_{R}$ unit via Eq. (2.5). To maintain compatibility with traditional information-theoretic expressions, the main text will primarily use the dimensionless form $D/\log\Gamma_{R}$; readers may directly interpret this as a physical quantity in units of $\mathrm{st}_{R}$.A minimal example: Consider a platform with $\Gamma_{R}=12$. If an experimental evaluation yields a relative entropy $D(\rho\|\sigma_{R})=\ln(144)$, the leading term of the complexity lower bound is calculated as $CR(\rho)=D_{\mathrm{nats}}/\ln(\Gamma_{R})=\ln(144)/\ln(12)=2~\mathrm{st}_{R}$. Ignoring logarithmic and spectral corrections, one can report a lower bound on the optimal circuit complexity as $C_{\mathrm{opt}}(\rho)\gtrsim 2~\mathrm{st}_{R}$. A standardized report for the leading term itself would be: $CR(\rho)=2~\mathrm{st}_{R}\;[\Gamma_{R}=12]$. To ensure clarity and reproducibility, we recommend following a minimal set of guidelines when reporting results using $\mathrm{st}_{R}$:
@@ 6.3 Resolving the Paradox of Pure State Complexity through an Observational Window
Proof Outline. Let $\sigma=E_{\Xi}(\rho)$. Apply a pinching map $\Delta_{P}(\sigma)=\sum_{i}P_{i}\sigma P_{i}$ to $\sigma$. By the entropy-increasing property of pinching, $S(\sigma)\leq S(\Delta_{P}(\sigma))$. At the same time, the entropy of the pinched state satisfies the block-entropy decomposition $S(\Delta_{P}(\sigma))=H(\{p_{\Xi,i}\})+\sum_{i}p_{\Xi,i}S(\sigma_{i})\leq H(\{p_{\Xi,i}\})+\sum_{i}p_{\Xi,i}\ln r_{i}$, where $\sigma_{i}=P_{i}\sigma P_{i}/p_{\Xi,i}$. Substituting this chain of inequalities back into the definition of $C_{R,\Xi}(\rho)$, $(\ln d_{R}-S(\sigma))/\ln\Gamma_{R}$, yields Eq. (6.2). (For technical details on the entropy increase due to pinching and the block-entropy decomposition used in this proposition, see Ref. [22]; for the rank-one dephasing entropy bound and experimental protocol, see D; for the axiomatic background of windowed monotonicity and notational/unit conventions, see Appendix I.) The physical insight of Eq. (6.2) is particularly crucial: it provides an operational "lever" to break the complexity degeneracy of pure states under a finite observational capability. For any pure state, the upper bound of the leading term of its ideal global complexity $C_{R}(\rho)$ is $\ln(d_{R})/\ln(\Gamma_{R})$. However, through a fixed local observational window $E_{\Xi}$, the value of $S(E_{\Xi}(\rho))$ for different pure states will differ significantly due to their different entanglement structures, thus creating a gap in their $C_{R,\Xi}(\rho)$ values. The significance of Eq. (6.3) is that it transforms $C_{R,\Xi}(\rho)$ from a theoretical construct into a physical quantity with direct operational meaning. While the definition of $C_{R,\Xi}(\rho)$ itself already conceptually solves the pure state degeneracy problem through $S(E_{\Xi}(\rho))$, Eq. (6.3) provides, from a more practical direction, an operational lower bound determined entirely by the classical measurement outcomes $\{p_{\Xi,i}\}$. This forms a key bridge connecting abstract theory with laboratory reality and gives rise to a powerful scheme for actively probing the latent complexity within pure states, which we term the "Experimental Squeeze of the Complexity Spectrum." Its core logic is that the choice of the measurement basis $\{P_{i}\}$ reflects the experimenter’s "informed probing" of the system’s intrinsic structural complexity. For a highly structured pure state, such as a graph state defined by a stabilizer group $S$, if we "intelligently" choose a measurement basis that is "aligned" with its stabilizer generators, then according to stabilizer theory, the measurement outcomes will be highly deterministic ($\pm 1$). This will cause the probability distribution $\{p_{\Xi,i}\}$ to become extremely sparse, with its Shannon entropy $H$ approaching zero. According to Eq. (6.3), this will cause its measurable complexity lower bound $C_{R,\Xi}(\rho)$ to be "squeezed" upwards, approaching its theoretical upper limit $\ln(d_{R})/\ln(\Gamma_{R})$. Conversely, a simple product state, lacking intrinsic non-local correlations, will inevitably appear more random when probed with the same complex measurement basis, resulting in a larger $H$ and a lower bound. This sharp contrast reveals a profound principle: the structural complexity of a state can only be most fully revealed when it is probed by an observer who "understands" its structural complexity. The significance of this "experimental squeeze" scheme is twofold. First, it constructs a feasible verification path for testing the core open question of this paper, Conjecture 6.1 (the near-tightness of RCC). By finding the optimal measurement bases that maximize the complexity lower bound for different classes of pure states (such as graph states, tensor network states, etc.), we can systematically establish "certificates of tightness" for their complexity. Second, it elevates the RCC framework from a static spectral measure to a complete, observation-dependent dynamical theoretical framework, in which the operational process of "finding the optimal measurement basis" is itself equivalent to the core task of "decoding the structural complexity of the quantum state."
6.4 A Dual-Constraint Framework: RCC and the Exact Quantum Speed Limit
The core contribution of this work is the establishment of a rigorous information-theoretic lower bound (Theorem 3.1) for any quantum process that prepares a given state, derived from the structured information of the final state. This bound is essentially a static constraint, uniquely determined by the information content of the final state $\rho$ relative to its structured vacuum $\sigma_{R}$, and is independent of the specific evolution path. A natural question follows: does a complementary constraint exist, one that originates from the process dynamics itself, to limit the efficiency of the evolution? Quantum Speed Limits (QSLs) provide the theoretical framework for this. Recent advancements, particularly the "exact-QSL," have established an upper bound on this speed, determined by the variance of the non-classical part of the generator relative to the instantaneous state—a bound that is tighter than the Mandelstam–Tamm bound [43] [44]. In this section, we juxtapose the RCC information-theoretic lower bound with the exact-QSL dynamical upper bound to construct a dual-constraint theoretical framework for quantum processes, revealing the complementary relationship between information cost and dynamical efficiency. We first consider a pure-state unitary evolution process driven by a Hamiltonian $H$, evolving from an initial state $|\psi_{0}\rangle$ to a final state $\rho_{T}:=|\psi_{T}\rangle\langle\psi_{T}|$. Following Hall’s exact uncertainty relation [45], the generator can be decomposed relative to the instantaneous state $\rho_{t}$ as $H=H_{\text{cl}}+H_{\text{nc}}$, where only the uncertainty of the non-classical part, $\Delta H_{\text{nc}}$, contributes to the instantaneous speed of the quantum state ray [44]. The energy uncertainty of an operator $X$ is defined as $\Delta X(t)=\sqrt{\langle X^{2}\rangle_{t}-\langle X\rangle_{t}^{2}}$, and the time average is defined as $\langle f\rangle_{T}:=T^{-1}\int_{0}^{T}[f(t)]\,dt$.
Definition 6.4.1 (Fubini-Study Angle). For an initial state $|\psi_{0}\rangle$ and a final state $|\psi_{T}\rangle$, the Fubini-Study angle $\theta$ between them is defined as[44]:
$\theta=\arccos(|\langle\psi_{0}|\psi_{T}\rangle|)\in[0,\pi/2]$ (6.4)
$\theta$ quantifies the shortest geometric distance between the two quantum state rays. The geometric distance that any quantum evolution process can "travel" in the state space is strictly limited by the cumulative effect of the Hamiltonian’s "non-classical driving force" throughout the process.
Theorem 6.4.1 (exact-QSL, adapted from [44]). For any pure-state unitary evolution, its Fubini-Study angle $\theta$ is constrained by the following dynamical upper bound:
$\theta\leq\frac{1}{\hbar}\int_{0}^{T}[\Delta H_{\text{nc}}(t)]\,dt\leq\frac{1}{\hbar}\int_{0}^{T}[\Delta H(t)]\,dt$ (6.5)
For a general dimension, an exact expression for the evolution time $T$ is:
$T=\hbar\frac{l(\phi_{t})|_{0}^{T}}{\langle\Delta H_{\text{nc}}\rangle_{T}}$ (6.6)
where $l(\phi_{t})$ is the Wootters/statistical distance [46], and $\theta\leq l(\phi_{t})|_{0}^{T}$. In cases such as an effective two-level system where the survival probability $p_{t}=|\langle\psi_{0}|\psi_{t}\rangle|^{2}$ is monotonic, or when the Hamiltonian $H^{2}=I$ (self-inverse), the equality $\theta=(1/\hbar)\int_{0}^{T}[\Delta H_{\text{nc}}(t)]\,dt$ holds. [^4] For a single-qubit system with a self-inverse Hamiltonian $H=\sigma_{x}$ (taking $\hbar=1$) and initial state $|\psi_{0}\rangle=|0\rangle$, the evolved state is $|\psi_{t}\rangle=\cos(t)|0\rangle-i\sin(t)|1\rangle$. Here, $\theta(t)=t$ and $\Delta H_{\text{nc}}(t)=1$, so the equality in (6.5) holds, providing a minimal, achievable example of this theory. We can now juxtapose this dynamical ceiling with the RCC information-theoretic floor from this work. The two bounds apply to different objects (circuit cost vs. state geometric distance) and do not form a sandwich bound on the same quantity.
Information-Theoretic Floor:
$C_{\text{opt}}^{(\epsilon)}(\rho_{T})\geq\frac{D(\rho_{T}\|\sigma_{R})}{\log(\Gamma_{R})}+\dots$ (6.7)
where $\dots"$ includes the universal logarithmic correction and the spectral skewness term $\Delta_{\text{spec}}$ (see Theorem 3.1); the dimension is normalized by $\log(\Gamma_{R})$ (in units of $st_{R}=\log_{2}(\Gamma_{R})$), consistent with the logarithmic base of $D(\cdot\|\cdot)$.
Dynamical Ceiling:
$\theta\leq\frac{1}{\hbar}\int_{0}^{T}[\Delta H_{\text{nc}}(t)]\,dt$ (6.8)
This framework reveals a physical picture: the total computational cost ($C_{\text{opt}}$) of any successful quantum preparation process must be high enough to surpass the "information-theoretic threshold" set by the final state’s information structure, while, simultaneously, its evolutionary efficiency at any moment is constrained by a "dynamical speed limit" determined by the non-classical part of the Hamiltonian. To explore the connection between these two on a heuristic level, we can introduce a complexity proxy based on geometric distance.
Definition 6.4.2 (State Geometric Complexity). For a given initial state $|\psi_{0}\rangle$, we define the state geometric complexity $C_{\text{FS}}$ of the final state $\rho_{T}$ as its Fubini-Study angle:
$C_{\text{FS}}(\rho_{T})=\theta$ (6.9)
A state with higher information-structural complexity (large RCC) typically implies that it contains more complex quantum correlations and a more refined distribution of probability amplitudes. Starting from a simple initial state (such as a vacuum or product state), constructing these complex structures intuitively requires the quantum state to traverse a longer, non-trivial geometric path in Hilbert space. This physical intuition inspires us that a positive correlation may exist between the state’s information-theoretic complexity (RCC) and its geometric complexity (quantified by the proxy $C_{\text{FS}}=\theta$). We therefore propose the following heuristic alignment:
$\text{RCC}_{R}(\rho_{T})\lesssim C_{\text{FS}}(\rho_{T})=\theta$ (6.10)
The $\lesssim$ symbol in Eq.(6.10) emphasizes that this is a heuristic alignment, not a universal proposition. This relation suggests that a state with higher information-structural complexity (large RCC) may also need to traverse a greater geometric distance to be reached from a simple initial state. It is crucial to contextualize this heuristic. In other frameworks, such as Nielsen’s geometric complexity which employs a Finsler metric with penalty factors, the relationship is often reversed, yielding a model-dependent lower bound of the form $C_{\text{Nielsen}}\geq\lambda_{\gamma}\theta$ [47]. Our heuristic alignment, therefore, specifically pertains to the comparison between the information content (RCC) and the minimal geometric path ($\theta$). With this understanding, we adopt a conservative formulation for experimental purposes, proposing the hypothesis $\theta\geq c_{h}\,RCC_{R}(\rho_{T})$, where $0<c_{h}\leq 1$ is a model-dependent heuristic constant. Combining this conservative hypothesis with the traditional QSL bound $T\geq\hbar\theta/\langle\Delta H\rangle_{T}$ leads to a meaningful physical conclusion.
Corollary 6.4.1 (Auditable Minimum Time Bound, Heuristic). If there exists a heuristic constant $c_{h}\in(0,1]$ such that $\theta\geq c_{h}\cdot\text{RCC}_{R}(\rho_{T})$, then the evolution time $T$ is constrained by the following lower bound, determined by the final state’s information structure:
$T\geq\frac{\hbar}{\langle\Delta H\rangle_{T}}\cdot\theta\geq\frac{\hbar c_{h}}{\langle\Delta H\rangle_{T}}\cdot\left[\frac{D(\rho_{T}\|\sigma_{R})}{\log(\Gamma_{R})}+\dots\right]$ (6.11)
This result connects the final state’s RCC information threshold with the process’s time-averaged energy uncertainty $\langle\Delta H\rangle_{T}$, collectively providing an auditable lower bound on the minimum evolution time. The dual-constraint framework proposed in this section unifies the static constraints of information theory and the dynamic constraints of dynamics, offering a more comprehensive and profound perspective for evaluating the resource consumption and evolutionary efficiency of quantum processes. It suggests that a deeper intrinsic unity may exist between computation, information, and dynamics than previously thought. Future open questions include: on which families of physically realizable systems does a stable quantitative relationship exist between RCC and $\theta$, and how to consistently generalize this framework to open quantum systems [48, 49].